OpenStudy (anonymous):

Help me with this please: -1/4(x-8)+1/2(x+2)=x+3

OpenStudy (radar):

do the parens getting -x/4+2+x/2+1=x+3 Multiply all terms by 4 clearing fractions. -x+8+2x+4=4x+12 consolidate similar terms: x-4x=12-12=0 -3x=0 or x=0

OpenStudy (radar):

As always substitute 0 for x in the original equation to verify.